# Measuring quantum relative entropy with finite-size effect

### v2

@@ C.2 Proof of Theorem 4
For $\alpha\in(0,1)$, we have
@@
$\leq$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
@@
Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(c)$ follows from (67).
@@
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $\mathrm{Tr}(\rho^{\otimes n})^{1+\alpha}(\sigma^{\otimes n})^{-\alpha}e^{-n\alpha(R-r)}+d_{n,d}e^{-nr}$
@@
$(d)$ follows from the monotonicity of sandwitch Rényi relative entropy.

### v4

@@ C.2 Proof of Theorem 4
Here, we define the projection $\{X>Y\}$ as the projection to the subspace spanned by the eigenvectors of $X-Y$ with positive eigenvalues. For $\alpha\in(0,1)$, we have
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}\kappa_{\sigma,n}(\sigma^{\otimes n})^{\alpha}e^{n\alpha R}$
@@
$\stackrel{{\scriptstyle(d)}}{{\leq}}$ $d_{n,d}^{\alpha}(\mathrm{Tr}(\sigma^{\frac{\alpha}{2(1-\alpha)}}\rho\sigma^{\frac{\alpha}{2(1-\alpha)}})^{1-\alpha}e^{\alpha R})^{n}$
@@
Here, $(b)$ follows from $\kappa_{\sigma,n}(\rho^{\otimes n})\leq I$, and $(d)$ follows from (70). The relation $(c)$ follows from
$\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{\kappa_{\sigma,n}(\sigma^{\otimes n})>e^{-nR}\kappa_{\sigma,n}(\rho^{\otimes n})\}$
$=$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}$
$\stackrel{{\scriptstyle(e)}}{{\leq}}$ $\mathrm{Tr}\kappa_{\sigma,n}(\rho^{\otimes n})^{1-\alpha}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}e^{n\alpha R}$
$\cdot\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}$
[…]
where $(e)$ follows from the following relation.
$\mathrm{Tr}$ $\Big{[}(\kappa_{\sigma,n}(\rho^{\otimes n}))^{1-\alpha}$
$\cdot\Big{(}(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}-(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}e^{n\alpha R}\Big{)}$
$\cdot\{e^{\alpha nR}(\kappa_{\sigma,n}(\sigma^{\otimes n}))^{\alpha}>(\kappa_{\sigma,n}(\rho^{\otimes n}))^{\alpha}\}\Big{]}\leq 0.$ (73)
@@
$\stackrel{{\scriptstyle(f)}}{{\leq}}$ $\mathrm{Tr}((\sigma^{\otimes n})^{-\frac{\alpha}{2(1+\alpha)}\rho^{\otimes n}(\sigma^{\otimes n})^{-\frac{\alpha}{2(1+\alpha)}})^{1+\alpha}e^{-\alpha(R-r)n}}$
$+d_{n,d}e^{-nr}$
@@
where $(f)$ follows from the monotonicity of sandwitch Rényi relative entropy.